The distance between the carbon and oxygen atoms in a CO molecule is 1.13 X 10-10m. Carbon has mass 12, and oxygen mass 16. Find the moment of inertia of the molecule if it rotates about an axis perpendicular to the line between the molecules and half way between them.
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..= 16 amu мо mų = 12 amu, mc mo - /ام /2 d d=1-13x10 m I= moment of inertia = Mc (2) + Mo. Mo (9) 16 +12) x 1067xTo nxcry]( 18